# Memory-augmented Transformers can implement Linear First-Order Optimization Methods

### v1

@@ Abstract
We show that memory-augmented Transformers (Memformers) can implement linear first-order optimization methods such as conjugate gradient descent, momentum methods, and more generally, methods that linearly combine past gradients. Building on prior work that demonstrates how Transformers can simulate preconditioned gradient descent, we provide theoretical and empirical evidence that Memformers can learn more advanced optimization algorithms. Specifically, we analyze how memory registers in Memformers store suitable intermediate attention values allowing them to implement algorithms such as conjugate gradient. Our results show that Memformers can efficiently learn these methods by training on random linear regression tasks, even learning methods that outperform conjugate gradient. This work extends our knowledge about the algorithmic capabilities of Transformers, showing how they can learn complex optimization methods.
@@ 1 Introduction
In-context learning (ICL) allows large language models (LLMs) to generate contextually appropriate outputs based solely on examples and queries provided in a prompt, without requiring any parameter adjustments (Brown, 2020; Liu et al., 2021; Lu et al., 2021; Wei et al., 2022; Wu et al., 2022). This remarkable ability has spurred research into understanding how Transformers can implement algorithms (Brown, 2020; Achiam et al., 2023; Touvron et al., 2023), with recent studies focusing on their capability to simulate optimization algorithms (Dai et al., 2022; Von Oswald et al., 2023a; Garg et al., 2022; Akyürek et al., 2022). Transformers have been shown to implement gradient-based optimization during their forward pass, such as preconditioned gradient descent for linear regression tasks (Dai et al., 2022; Mahankali et al., 2023; Ahn et al., 2024).
More recently, studies have demonstrated that Transformers can learn even more advanced optimization methods. For instance, Fu et al. (2023) showed that Transformers exhibit convergence rates comparable to Iterative Newton’s Method, a higher-order optimization technique that converges exponentially faster than gradient descent for in-context linear regression. Additionally, Vladymyrov et al. (2024) proved that Transformers can, in fact, learn a variant of gradient descent that approximates second-order methods, such as $\mathrm{GD}^{++}$, achieving convergence rates similar to Newton’s method. These findings lead to the central question of our paper:
@@
We hope that by addressing this question, we can reveal some of the representational power of Transformers as “algorithm learners,” and thus further motivate the use of machine learning for discovering new optimization algorithms. To make our question more precise, we focus on learning the entire class of gradient-based algorithms obtained by linearly combining past gradients, known as Linear First-Order Methods (LFOMs) (Goh, 2017), where the $k+1^{\text{st}}$ iterate is
$w^{k+1}=w^{0}+\sum_{i=0}^{k}\Gamma_{i}^{k}\nabla f(w^{i}),$ (1.1)
@@
and where $\{\Gamma_{i}^{k}\}_{i=0}^{k}$ are diagonal matrices. Model (1.1) is quite general, as it includes, as special cases, standard methods such as gradient descent (GD), momentum GD, Nesterov’s accelerated gradient, conjugate gradient, and in a stochastic setting, AdaGrad, ADAM, among others.
@@
Our key insight for efficiently learning LFOMs is to leverage memory-augmented Transformers, known as Memformers (Wu et al., 2020; Xu et al., 2021), which retain intermediate attention values across layers. This memory can enable them to store past gradients, facilitating the execution of advanced first-order methods such as conjugate gradient descent and momentum methods. The same mechanism allows Memformers to implement more general LFOMs.
@@
While unconditional learning of gradient methods remains out of reach, we build on related work demonstrating that Transformers can learn gradient descent in the context of linear regression tasks (Garg et al., 2022; Akyürek et al., 2022; Von Oswald et al., 2023a; Ahn et al., 2024; Zhang et al., 2024). Inspired by these findings, and extending the work of (Ahn et al., 2024), we conduct a theoretical analysis of the loss landscape for memory-augmented linear Transformers that omit softmax activation (Schlag et al., 2021; Von Oswald et al., 2023a; Ahn et al., 2024).
In summary, our main contributions are as follows:
Main Contributions
  1. (A)
    Theoretical and Experimental Evidence. We provide theoretical justification and empirical results showing that Memformers trained on linear regression tasks learn LFOMs (and even quasi-Newton methods like $\mathrm{GD}^{++}$ (Von Oswald et al., 2023a) for novel in-context data.
  2. (B)
    Performance Beyond Conjugate Gradient Descent. We demonstrate that LFOMs learned by Memformers significantly outperform conjugate gradient descent on training data, while remaining competitive on test data, indicating good generalization performance.
  3. (C)
    Multi-Headed Attention Improves Performance. We show empirically that multi-headed attention improves Memformers’ test performance, and offer theoretical justification for why increasing attention heads enhances loss performance on test data.
The main objective of this paper is to investigate the potential of augmented Transformers to learn optimization algorithms. We are not advocating for Transformers as substitutes for established optimization methods, but rather, we aim to kindle further exploration into their algorithmic capabilities. We believe our results contribute to understanding how augmented Transformers could help with optimization, which may ultimately lead to new and practical gradient-based algorithms.
@@ 1.1 Related Work
Research on Transformers is extremely active, and we cannot hope to fully capture the breadth of the related literature. Below, we summarize the most immediately relevant topics.
In-Context Learning. The ability of Transformer models to perform in-context learning (ICL) has been extensively studied since its introduction by Brown (2020). Subsequent works have explored how these models adapt to new tasks without requiring parameter updates (Xie et al., 2021; Von Oswald et al., 2023b; Hahn and Goyal, 2023; Liu et al., 2021; Lu et al., 2021; Wei et al., 2022; Wu et al., 2022). This foundational research has paved the way for studies investigating how Transformers can implement specific algorithms, such as gradient-based methods.
Gradient-Based Methods in Transformers. Garg et al. (2022) analyze the learning of gradient descent within Transformers, particularly in the context of ICL for linear functions. Empirical studies (Garg et al., 2022; Akyürek et al., 2022; Von Oswald et al., 2023a) have shown that Transformers can learn gradient descent after being trained on random linear regression tasks. Expanding on these results, Von Oswald et al. (2023a); Ahn et al. (2024) demonstrate that Transformers can implement preconditioned gradient descent for solving linear regression problems presented in input prompts. Notably, these works—as well as ours—utilize Linear Transformers as discussed in (Schlag et al., 2021; Von Oswald et al., 2023a; Ahn et al., 2023).
@@
Memory-Augmented Transformers (Memformers). Memformers were introduced by Wu et al. (2020); Xu et al. (2021). These models retain intermediate attention values across layers through memory registers, enabling more complex computations and optimization methods to be learned. While significant progress has been made in understanding how Transformers can learn gradient descent, their potential for implementing more sophisticated LFOMs remains largely unexplored. Our work addresses this gap by showing how Memformers can efficiently implement a wide range of advanced first-order and quasi-second-order optimization techniques, including CGD and momentum methods, thereby pushing the boundaries of Transformer-based architectures.
@@ 2.1 Linear Transformers on Random Linear Regression
We follow the setup of training Transformers on random instances of linear regression, following the prior works (Garg et al., 2022; Akyürek et al., 2022; Von Oswald et al., 2023a; Ahn et al., 2024). We largely use the notation and formal setup of (Ahn et al., 2024), which we now proceed to recall.
Data Distribution. Let $\mathbf{x}(i)\in\mathbb{R}^{d}$ represent covariates drawn independently from a distribution $\mathcal{D}_{\mathbf{X}}$, and let $\mathbf{w}^{*}\in\mathbb{R}^{d}$ be drawn from $\mathcal{D}_{\mathbf{W}}$. The matrix of covariates $\mathbf{X}\in\mathbb{R}^{(n+1)\times d}$ contains rows $\mathbf{x}(i)$. The responses are $\mathbf{y}=[\langle\mathbf{x}(1),\mathbf{w}^{*}\rangle,\dots,\langle\mathbf{x}(n),\mathbf{w}^{*}\rangle]\in\mathbb{R}^{n}$. Define the input matrix $\mathbf{Z}_{0}$ as:
$\mathbf{Z}_{0}=\begin{bmatrix}\mathbf{x}(1)&\mathbf{x}(2)&\cdots&\mathbf{x}(n)&\mathbf{x}(n+1)\\ \mathbf{y}(1)&\mathbf{y}(2)&\cdots&\mathbf{y}(n)&0\end{bmatrix}\in\mathbb{R}^{(d+1)\times(n+1)},$ (2.1)
@@
Transformer Architecture. As in the related work, we also simplify the Transformer to consider only attention layers, using $L$ layers of linear self-attention with a residual connection. Therefore, for each layer $\ell$, the output is updated as
@@
Using updates (2.5), with the input $\mathbf{Z}_{0}$, the final transformer output is
@@
$f\left(\{P_{\ell},Q_{\ell}\}_{\ell=0}^{L-1}\right)=\mathbb{E}_{(\mathbf{Z}_{0},\mathbf{w}^{*})}\left[\left(\text{TF}_{L}(\mathbf{Z}_{0})+(\mathbf{w}^{*})^{\top}\mathbf{x}(n+1)\right)^{2}\right].$ (2.7)
Here, the scaling factor $\frac{1}{n}$ is used only for ease of notation and does not influence the expressive power of the Transformer.
We will utilize the following lemma from Ahn et al. (2024), which demonstrates that multi-layer Transformers simulate preconditioned gradient descent under suitable parameterization. We have provided the full proof of this Lemma 2.11 in the Appendix for completeness.
$P_{\ell}=\begin{bmatrix}\mathbf{B}_{\ell}=0_{d\times d}&0\\ 0&1\end{bmatrix},\quad Q_{\ell}=-\begin{bmatrix}\mathbf{A}_{\ell}&0\\ 0&0\end{bmatrix},\quad\mathbf{A}_{\ell},\mathbf{B}_{\ell}\in\mathbb{R}^{d\times d}.$ (2.8)
Lemma 1 (Lemma 1, Ahn et al. (2024)).
Consider an $L$-layer linear transformer parameterized by $\mathbf{A}_{0},\dots,\mathbf{A}_{L-1}$, as in (2.8). Let $y_{\ell}^{(n+1)}$ be the $(d+1,n+1)$-th entry of the $\ell$-th layer output, i.e., $y_{\ell}^{(n+1)}=[\mathbf{Z}_{\ell}]_{(d+1),(n+1)}$ for $\ell=1,\dots,L$.
@@
$R_{\mathbf{w}^{*}}(\mathbf{w}):=\frac{1}{2n}\|\mathbf{X}^{\top}\mathbf{w}-\mathbf{X}^{\top}\mathbf{w}^{*}\|^{2}=\frac{1}{2n}(\mathbf{w}-\mathbf{w}^{*})^{\top}\mathbf{X}\mathbf{X}^{\top}(\mathbf{w}-\mathbf{w}^{*}).$ (2.11)
Note: The Hessian of $R_{\mathbf{w}^{*}}(\mathbf{w})$ is given by:
$\nabla^{2}R_{\mathbf{w}^{*}}(\mathbf{w})=\frac{1}{n}\mathbf{X}\mathbf{X}^{\top}.$ (2.12)
2.2 Linear First-Order Methods
Linear First-Order Methods (LFOMs) (Goh, 2017) are a class of optimization algorithms that lineary combine past gradients for minimizing smooth objective functions. They iteratively update a parameter vector $\mathbf{w}$ using the gradient of the objective function. The general update rule is
@@ 2.2 Linear First-Order Methods
where $\alpha_{k}$ is the step size and $\mathbf{d}^{k}$ is the update direction, typically related to the gradient $\nabla f(\mathbf{w}^{k})$. Algorithms within this family differ in how they compute $\mathbf{d}^{k}$ and choose $\alpha_{k}$.
LFOMs can be expressed in a cumulative form. For gradient descent, unrolling (2.13) we get
$\mathbf{w}^{k+1}=\mathbf{w}^{0}-\alpha\sum\nolimits_{i=0}^{k}\nabla f(\mathbf{w}^{i}),$ (2.14)
while common momentum methods need an additional term incorporating past gradients, yielding
$\mathbf{w}^{k+1}=\mathbf{w}^{0}+\sum\nolimits_{i=0}^{k}\gamma_{i}^{k}\nabla f(\mathbf{w}^{i}),$ (2.15)
where the coefficients $\gamma_{i}^{k}$ weight previous gradients. More advanced methods, or general LFOMs, use diagonal matrices $\Gamma_{i}^{k}$ to coordinate-wise scale each gradient component, i.e.,
$\mathbf{w}^{k+1}=\mathbf{w}^{0}+\sum\nolimits_{i=0}^{k}\Gamma_{i}^{k}\nabla f(\mathbf{w}^{i}).$ (2.16)
Momentum Methods and Conjugate Gradient Descent (CGD)
Momentum methods accelerate convergence by incorporating a momentum term, modifying the gradient to account for past updates and achieving faster convergence in relevant directions. Conjugate Gradient Descent (CGD), on the other hand, is a first-order method optimized for quadratic minimization, serving as a benchmark for large-scale, sparse linear systems. After an initial steepest descent, CGD generates directions conjugate to previous ones, leading to faster convergence than standard gradient descent. Both are core methods within the LFOM class, summarized below:
Momentum Methods
1:Initialize $\mathbf{w}_{0}$, $\mathbf{v}_{0}=0$
2:for $n=1,2,\dots$ do
3:     Compute the gradient:
$\nabla f(\mathbf{w}_{n})$
4:     Update the velocity:
$\mathbf{v}_{n}=\beta\mathbf{v}_{n-1}-\eta\nabla f(\mathbf{w}_{n})$
5:     Update the iterate:
$\mathbf{w}_{n+1}=\mathbf{w}_{n}+\mathbf{v}_{n}$
6:end for
7:$\beta$: Momentum coefficient (controls the influence of past gradients)
8:$\eta$: Learning rate (scales the gradient step size)
Conjugate Gradient Descent (CGD)
1:Initialize $\mathbf{w}_{0}$, $\mathbf{s}_{0}=-\nabla f(\mathbf{w}_{0})$
2:for $n=1,2,\dots$ do
3:     Compute the steepest descent direction:
$\Delta\mathbf{w}_{n}=-\nabla f(\mathbf{w}_{n})$
4:     Compute the conjugacy coefficient:
$\gamma_{n}=\frac{\|\nabla f(\mathbf{w}_{n})\|^{2}}{\|\nabla f(\mathbf{w}_{n-1})\|^{2}}$
5:     Update the search direction:
$\mathbf{s}_{n}=\Delta\mathbf{w}_{n}+\gamma_{n}\mathbf{s}_{n-1}$
6:     Perform a line search:
$\alpha_{n}=\arg\min_{\alpha}f(\mathbf{w}_{n}+\alpha\mathbf{s}_{n})$
7:     Update the iterate:
$\mathbf{w}_{n+1}=\mathbf{w}_{n}+\alpha_{n}\mathbf{s}_{n}$
8:end for
Momentum methods provide fast convergence by accumulating gradient history and are widely used in modern optimization. CGD converges in at most $N$ iterations for quadratic functions, where $N$ is the number of variables, and is effective for ill-conditioned problems.
3 Memformers Implement LFOMs In-Context
Building on Lemma 2.11, we demonstrate how memory-augmented Transformers can implement Linear First-Order Methods (LFOMs) like Conjugate Gradient Descent (CGD) in-context. As noted in Ahn et al. (2024) (Subsection C.1), the term $\text{Attn}_{P_{\ell},Q_{\ell}}(\mathbf{Z}_{\ell})$ in the update for $\mathbf{Z}_{\ell+1}$ corresponds to the preconditioned gradient $\mathbf{A}_{\ell}\nabla R_{\mathbf{w}^{*}}(\mathbf{w}_{\ell}^{\mathrm{gd}})$ of the in-context loss (2.11).
3.1 Dynamic Memory for CGD-like Algorithms
With attention simulating preconditioned gradient descent, we now explore how memory-augmented Transformers can simulate Conjugate Gradient Descent (CGD). The dynamic memory mechanism enables recursive refinement of search directions, closely mimicking CGD’s iterative process.
Proposition 1.

### v3

# Toward generalizable learning of all (linear) first-order methods via memory augmented Transformers

@@ Abstract
We show that memory-augmented Transformers can implement the entire class of linear first-order methods (LFOMs), a class that contains gradient descent (GD) and more advanced methods such as conjugate gradient descent (CGD), momentum methods and all other variants that linearly combine past gradients. Building on prior work that studies how Transformers simulate GD, we provide theoretical and empirical evidence that memory-augmented Transformers can learn more advanced algorithms. We then take a first step toward turning the learned algorithms into actually usable methods by developing a mixture-of-experts (MoE) approach for test-time adaptation to out-of-distribution (OOD) samples. Lastly, we show that LFOMs can themselves be treated as learnable algorithms, whose parameters can be learned from data to attain strong performance.
@@ 1 Introduction
In-context learning (ICL) allows large language models (LLMs) to generate contextually appropriate outputs based solely on examples and queries provided in a prompt (Brown, 2020; Liu et al., 2021; Lu et al., 2021; Wei et al., 2022; Wu et al., 2022). This remarkable ability has spurred research into understanding how Transformers can implement algorithms (Achiam et al., 2023; Touvron et al., 2023), with recent studies focusing on they can simulate optimization algorithms (Dai et al., 2022; Von Oswald et al., 2023a; Garg et al., 2022; Akyürek et al., 2022). Transformers have been shown to implement gradient-based optimization during their forward pass, such as preconditioned gradient descent for linear regression tasks (Dai et al., 2022; Mahankali et al., 2023; Ahn et al., 2024).
Other recent work shows that Transformers can also learn more advanced optimization methods, e.g., Fu et al. (2023) show that Transformers exhibit convergence rates comparable to Newton’s Method for ICL on linear regression. Vladymyrov et al. (2024) prove that Transformers can learn a variant of gradient descent that approximates second-order methods, such as $\mathrm{GD}^{++}$, achieving convergence rates similar to Newton’s method. These findings lead to the central question of our paper:
@@
We address this question by investigating the representational power of Transformers as algorithm learners, and with it, contributing to the topic of machine driven algorithm discovery. More specifically, we focus on learning all gradient-based algorithms obtainable by linearly combining past gradients, known as Linear First-Order Methods (LFOMs) (Goh, 2017), wherein $(k+1)$-st iterate is
$w^{k+1}=w^{0}+\sum\nolimits_{i=0}^{k}\Gamma_{i}^{k}\nabla f(w^{i}),$ (1.1)
@@
and where $\{\Gamma_{i}^{k}\}_{i=0}^{k}$ are diagonal matrices. Iteration (1.1) is quite general: it includes as special cases standard methods such as gradient descent (GD), momentum GD, Nesterov’s accelerated gradient, conjugate gradient, and in a stochastic version, AdaGrad, ADAM, among others.
@@
Toward showing how Transformers can (efficiently) capture all LFOMs, our key insight is to consider memory-augmented Transformers, known as Memformers (Wu et al., 2020; Xu et al., 2021), which retain intermediate attention values across layers. This memory enables Memformers to store past gradients, and facilitates efficient mimicking of first-order methods such as conjugate gradient descent and momentum methods, that also use past gradients.
One may ask: What does it mean to “learn” an optimization algorithm? For us, it refers to two key aspects:
1. Expressivity. The Memformer can perform iterations of LFOMs in its forward pass, under a suitable choice of parameters. Thus, the architecture and parameterization are sufficiently expressive to simulate LFOMs.
2. Trainability. The Memformer’s parameters can be trained on random linear regression tasks. Then, using these learned parameters (which are shared across all in-context data samples), the Memformer can execute “CGD-like” and “LFOM-like” iterations during a forward pass. (cf. Section 3)
@@
While far from unconditional learning of GD, the above two aspects of learning follow a large body of work on how Transformers learn GD in context (Garg et al., 2022; Akyürek et al., 2022; Von Oswald et al., 2023a; Ahn et al., 2024; Zhang et al., 2024). Inspired by these works, and extending the work of Ahn et al. (2024), we study “learning” by analyzing the loss landscape for memory-augmented Linear Transformers that omit softmax activations (Schlag et al., 2021; Von Oswald et al., 2023a; Ahn et al., 2024).
1.1 Main Contributions
  1. 1.
    Theoretical justification that Memformers can implement LFOM iterations, including CGD. We provide a rigorous theoretical framework showing that Memformers, when trained on linear regression tasks, can be configured to perform iterations of any desired LFOM in their forward pass.
  2. 2.
    Empirical evidence that Memformers learn optimization algorithms. Through extensive experiments, we demonstrate that a Memformer can learn LFOMs by training on random linear regression tasks. Thereafter, using its learned parameters, it can solve new regression problems using its forward pass. Notably, on several inputs such a Memformer performs competitively against (sometimes even outpeforming) CGD, Nesterov AGM, and momentum GD.
    This finding is potentially significant because while CGD (and several LFOMs) adapts its step-size and conjugacy parameters per sample—which translates into $2B$ parameters $\{\alpha,\gamma\}$ per iteration, for a batch of size $B$—a Memformer reuses one set of learned weights across the entire batch (see also Remark 3.3).
  3. 3.
    Enhanced performance through multi-headed attention and mixture-of-experts (MoE). We demonstrate that multi-headed attention improves both in-distribution and out-of-distribution (OOD) performance, with an MoE approach that enables each attention head to specialize in handling different data distributions. Additionally, we explore the performance of LFOMs when their parameters are learned directly from data, thus framing LFOMs as statistically learnable algorithms, and showcase their ability to generalize effectively across diverse in-context data samples.
  4. 4.
    Connections to test-time adaptation and relevance for LLMs. Beyond in-context adaptation, we investigate ways to update model parameters at test time to address shifting distributions, a topic with growing importance in large language models (LLMs). Recent work suggests that appropriately scaling test-time compute, rather than just model size, can yield stronger performance improvements (Snell et al., 2024), and we link these insights to our MoE-based Memformer approach (Section 4.2) to demonstrate how gating and memory preconditioners can adapt at inference time to handle various distribution shifts (Sun et al., 2020).
Our objective. We would like to emphasize that our main objective is to investigate the potential of Memformers to learn advanced optimization algorithms in a general sense, as well as their potential to adapt to OOD optimization problems. We are not advocating for Transformers as replacement for established optimization methods. Nevertheless, we hope that in the future, machine-discovered optimization algorithms surpass hand-designed methods such as GD or CGD, at least in an average-case / distributional sense.
@@ 1.2 Related Work
In-Context Learning. The ability of Transformer models to perform in-context learning (ICL) has been extensively studied since its introduction by Brown (2020). Subsequent works have explored how these models adapt to new tasks without requiring parameter updates (Xie et al., 2021; Von Oswald et al., 2023b; Hahn and Goyal, 2023; Liu et al., 2021; Lu et al., 2021; Wei et al., 2022; Wu et al., 2022).
Gradient-Based Methods in Transformers. Garg et al. (2022) analyze the learning of GD within Transformers, particularly in the context of ICL for linear functions. Empirical studies (Garg et al., 2022; Akyürek et al., 2022; Von Oswald et al., 2023a) have shown that Transformers can learn GD after being trained on random linear regression tasks. Expanding on these results, Von Oswald et al. (2023a); Ahn et al. (2024) demonstrate that Transformers can implement preconditioned GD for solving linear regression problems presented in input prompts. Notably, these works, as well as ours, utilize Linear Transformers as discussed in (Schlag et al., 2021; Von Oswald et al., 2023a; Ahn et al., 2023).
@@
Memory-Augmented Transformers (Memformers). Memformers were introduced in (Wu et al., 2020; Xu et al., 2021). While significant progress has been made in understanding how Transformers can learn GD, their potential for learning more sophisticated methods remains largely unexplored. Our work addresses this gap and shows how Memformers can efficiently implement a wide range of advanced first-order and quasi-second-order methods.
@@ 2 Background and Problem Setup
We follow the setup of training Transformers on random instances of linear regression, following the prior works (Garg et al., 2022; Akyürek et al., 2022; Von Oswald et al., 2023a; Ahn et al., 2024). We largely follow the notation and formal setup of (Ahn et al., 2024), which we recall below.
@@ 2.1 Linear Transformers on Random Linear Regression
Data Distribution. Let $\mathbf{x}(i)\in\mathbb{R}^{d}$ represent covariates drawn independently from a distribution $\mathcal{D}_{\mathbf{X}}$, and let $\mathbf{w}^{*}\in\mathbb{R}^{d}$ be drawn from $\mathcal{D}_{\mathbf{W}}$. The matrix of covariates $\mathbf{X}\in\mathbb{R}^{(n+1)\times d}$ contains rows $\mathbf{x}(i)$. The responses are $\mathbf{y}=[\langle\mathbf{x}(1),\mathbf{w}^{*}\rangle,\dots,\langle\mathbf{x}(n),\mathbf{w}^{*}\rangle]\in\mathbb{R}^{n}$. Define the input matrix $\mathbf{Z}_{0}\in\mathbb{R}^{(d+1)\times(n+1)}$ as:
$\mathbf{Z}_{0}=\begin{bmatrix}\mathbf{x}(1)&\mathbf{x}(2)&\cdots&\mathbf{x}(n)&\mathbf{x}(n+1)\\ \mathbf{y}(1)&\mathbf{y}(2)&\cdots&\mathbf{y}(n)&0\end{bmatrix},$ (2.1)
@@
Architecture. Following the related work, we also simplify the Transformer to consider only attention layers, using $L$ layers of linear self-attention with a residual connection. Therefore, for each layer $\ell$, the output is updated as
@@
Using (2.5), with the input $\mathbf{Z}_{0}$, the final Transformer output is
@@
$\mathbb{E}_{(\mathbf{Z}_{0},\mathbf{w}^{*})}\left[\left(\text{TF}_{L}(\mathbf{Z}_{0})+(\mathbf{w}^{*})^{\top}\mathbf{x}(n+1)\right)^{2}\right].$ (2.7)
We will utilize the following lemma from Ahn et al. (2024), which demonstrates that multi-layer Transformers simulate preconditioned gradient descent under suitable parameterization. We have provided the full proof of this lemma in Appendix A for completeness.
$P_{\ell}=\begin{bmatrix}\mathbf{B}_{\ell}=0_{d\times d}&0\\ 0&1\end{bmatrix},\quad Q_{\ell}=-\begin{bmatrix}\mathbf{A}_{\ell}&0\\ 0&0\end{bmatrix}$ (2.8)
Lemma 2.1 (Lemma 1, Ahn et al. (2024)).
Let an $L$-layer linear transformer be parameterized by $\mathbf{A}_{0},\dots,\mathbf{A}_{L-1}$, as in (2.8). Let $y_{\ell}^{(n+1)}=[\mathbf{Z}_{\ell}]_{(d+1),(n+1)}$ for $\ell=1,\dots,L$; then,
@@
$R_{\mathbf{w}^{*}}(\mathbf{w}):=\frac{1}{2n}\|\mathbf{X}^{\top}\mathbf{w}-\mathbf{X}^{\top}\mathbf{w}^{*}\|^{2}.$ (2.11)
Note. As observed in §C.1 of Ahn et al. (2024), the term $\mathrm{Attn}_{P_{\ell},Q_{\ell}}(\mathbf{Z}_{\ell})$ in (2.5) corresponds to the preconditioned gradient $\mathbf{A}_{\ell}\nabla R_{\mathbf{w}^{*}}(\mathbf{w}_{\ell}^{\mathrm{gd}})$ of (2.11) in the update for $\mathbf{w}_{\ell+1}^{\mathrm{gd}}$.
2.2 Linear First-Order Methods (LFOMs)
Linear First-Order Methods (LFOMs) (Goh, 2017) optimize smooth functions by iteratively updating a parameter vector $\mathbf{w}$ with current and past gradients. The general update is
@@ 2.2 Linear First-Order Methods (LFOMs)
where $\alpha_{k}$ is a step size and $\mathbf{d}^{k}$ is a direction typically tied to $\nabla f(\mathbf{w}^{k})$. Different LFOMs vary in how $\mathbf{d}^{k}$ and $\alpha_{k}$ are chosen. For example:
[…]
In more advanced LFOMs, the scalars $\gamma_{i}^{k}$ are replaced by diagonal matrices $\Gamma_{i}^{k}$, enabling coordinate-wise scaling.
Conjugate Gradient Descent (CGD). Specialized for quadratic minimization, CGD generates update directions $\mathbf{s}_{n}$ that are conjugate to previous ones, leading to faster convergence than standard gradient descent. It iterates:
$\Delta\mathbf{w}_{n}=-\nabla f(\mathbf{w}_{n}),\quad\gamma_{n}=\frac{\|\nabla f(\mathbf{w}_{n})\|^{2}}{\|\nabla f(\mathbf{w}_{n-1})\|^{2}},$ (2.13)
$\mathbf{s}_{n}=\Delta\mathbf{w}_{n}+\gamma_{n}\mathbf{s}_{n-1},$
then chooses a step size $\alpha_{n}$ (often via line search) and updates: $\mathbf{w}_{n+1}=\mathbf{w}_{n}+\alpha_{n}\,\mathbf{s}_{n}.$
Both momentum GD and CGD are LFOMs. Momentum methods are widely adopted in modern optimization, while CGD converges in at most $N$ iterations for $N$-dimensional quadratics, and is effective for ill-conditioned problems.
Figure 1: CGD-like Memformer (3.2) without preconditioning ($\mathbf{A}_{\ell}=\mathbf{I}$) vs. actual CGD running separately on each test sample. Test data is drawn from the same distribution as the training data.
Figure 2: CGD-like Memformer (3.2) with preconditioning ($\mathbf{A}_{\ell}\neq\mathbf{I}$). This yields a more general LFOM-like scheme, often outperforming CGD, Nesterov AGM and momentum GD. Test data is independently drawn from the same distribution as training data.
Figure 3: LFOM Memformer (3.4) vs. CGD, Nesterov AGM and momentum GD (Pre = non-trivial preconditioners).
3 Memformers Can Implement LFOMs
Memformers can “learn” LFOMs in the sense described in Section 1. Each layer $\ell$ has learnable parameters $\mathbf{A}_{\ell},\mathbf{B}_{\ell}$ (2.8) and either $\alpha_{\ell},\gamma_{\ell}$ (3.2) or $\Gamma_{\ell}$ (3.4). Theorems 3.1 and 3.2 show that with suitable parameters, Memformers can exactly implement CGD and LFOM iterations in their forward pass.
With only a small number of learned parameters shared across a batch of in-context test data samples, Memformers can perform “CGD-like” (Section 3.1) or “LFOM-like” (Section 3.2) updates that rival or even surpass CGD. We refer to algorithms realized by (3.2) as “CGD-like” and those realized by (3.4) as “LFOM-like”.
3.1 Single Dynamic Memory for CGD-like Algorithms
Theorem 3.1.
A memory-augmented Transformer can implement Conjugate Gradient Descent (CGD) in its forward pass via a single memory register, where: